### v1

# InternLM-XComposer2.5-Reward: A Simple Yet Effective Multi-Modal Reward Model

Abstract
Despite the promising performance of Large Vision Language Models (LVLMs) in visual understanding, they occasionally generate incorrect outputs. While reward models (RMs) with reinforcement learning or test-time scaling offer the potential for improving generation quality, a critical gap remains: publicly available multi-modal RMs for LVLMs are scarce, and the implementation details of proprietary models are often unclear. We bridge this gap with InternLM-XComposer2.5-Reward (IXC-2.5-Reward), a simple yet effective multi-modal reward model that aligns LVLMs with human preferences. To ensure the robustness and versatility of IXC-2.5-Reward, we set up a high-quality multi-modal preference corpus spanning text, image, and video inputs across diverse domains, such as instruction following, general understanding, text-rich documents, mathematical reasoning, and video understanding. IXC-2.5-Reward achieves excellent results on the latest multi-modal reward model benchmark and shows competitive performance on text-only reward model benchmarks. We further demonstrate three key applications of IXC-2.5-Reward: (1) Providing a supervisory signal for RL training. We integrate IXC-2.5-Reward with Proximal Policy Optimization (PPO) yields IXC-2.5-Chat, which shows consistent improvements in instruction following and multi-modal open-ended dialogue; (2) Selecting the best response from candidate responses for test-time scaling; and (3) Filtering outlier or noisy samples from existing image and video instruction tuning training data. To ensure reproducibility and facilitate further research, we have open-sourced all model weights and training recipes at https://github.com/InternLM/InternLM-XComposer.
[…]
1 Introduction
If you don’t know where you are going, you’ll end up some
place else.
— Yogi Berra
Figure 1: (a) To train the IXC-2.5-Reward, we construct a multi-modal preference dataset spanning diverse domains (e.g., natural scenes, text-rich, reasoning) and modalities (image, text, video). (b) The framework of IXC-2.5-Reward. (c) The IXC-2.5-Reward guides policy training for IXC-2.5-Chat via reinforcement learning.
Reward Models (RMs) [8, 121, 48, 92, 88, 104, 55, 98, 105, 78, 91] provide the crucial direction guidance about how well an AI model’s outputs align with human preference, and benefit Large Language Models (LLMs) in training and inference. During training, RMs are often used with reinforcement learning from human feedback (RLHF) [66, 7, 71, 70] to penalize undesirable model behaviors and encourage outputs that align with human values. At inference, RMs facilitate test-time scaling strategies [80, 27], such as selecting the best response from candidate outputs or providing step-by-step critiques for complex reasoning tasks [107, 29].
Despite their crucial role in both training and inference, multi-modal RMs for Large Vision Language Models (LVLMs) remain notably underexplored compared to language-only RMs for LLMs. Because current preference data is predominantly text-based and skewed toward specific domains (e.g., safety), data scarcity poses a significant challenge to training multi-modal RMs for diverse modalities such as images, videos, and text. Consequently, existing multi-modal RMs [87, 95] are largely constrained to narrow domains (e.g., mitigating hallucinations) or rely on prompting LVLMs with evaluation prompts, effectively functioning as generative RMs [94]. The limitation of multi-modal RMs subsequently constrains the capabilities of open-source LVLMs such as instruction following and safety-should-refuse, thereby hampering user interaction experience in multi-modal chat scenarios.
The growing community interest in RLHF and test-time scaling highlights the need for multi-modal RMs, which motivates us to present InternLM-XComposer2.5-Reward (IXC-2.5-Reward). Instead of directly transferring unimodal (text) reward models (RMs) to the vision modality, we augment the existing LVLM (InternLM-XComposer2.5) with an additional scoring head to predict reward scores. An effective multi-modal RM should ideally possess two key properties: (1) the ability to predict reward scores for both image, video, and textual inputs and (2) the capacity to generalize across diverse domains, such as instruction following, knowledge, text-rich images (e.g., documents), reasoning tasks, etc. To this end, we develop a pipeline (Fig. 1(a)) to construct multi-modal preference data, and also incorporate existing high-quality datasets. This pipeline selects prompts across diverse domains for text, image, and video inputs, generates corresponding responses, and then uses GPT-4o [31] or verifier [40] to perform preference judgments. Trained on our preference data, IXC-2.5-Reward effectively evaluates both visual (image and video) and textual inputs (Fig. 1 (b)).
IXC-2.5-Reward achieves best performance on multi-modal VL-RewardBench [43] (70.0%) that beat all previous generative RMs including Gemini-1.5-Pro (62.5%) and GPT-4o (62.4%). Even on uni-modal (text) RM benchmarks, IXC-2.5-Reward also demonstrates good results, with an average score of $88.6\%$ on Reward-Bench [41] and $68.8\%$ on RM-Bench [51].
We further demonstrate the effectiveness of IXC-2.5-Reward in the following three aspects:
(1) IXC-2.5-Reward for RL training. We train a chat model (IXC-2.5-Chat) using the on-policy Proximal Policy Optimization (PPO) algorithm with IXC-2.5-Reward to enhance its ability to follow instructions and provide a better user experience in multi-modal conversations. Our results show clear improvements of IXC-2.5-Chat on multi-modal instruction following and in-the-wild chatting benchmarks, which validate the effectiveness of IXC-2.5-Reward for providing the reward signal during RL training.
(2) IXC-2.5-Reward for Test-Time Scaling. Using best-of-$N$ sampling with IXC-2.5-Reward leads to additional performance gains compared to the RL-trained IXC-2.5-Chat, confirming IXC-2.5-Reward’s effectiveness in selecting good responses from candidate responses.
(3) IXC-2.5-Reward for Data Cleaning. We observe a strong correlation between low IXC-2.5-Reward scores and problematic samples, such as those exhibiting hallucinations or mismatched image/video and question/answer content. This suggests that IXC-2.5-Reward can effectively clean LVLM pre-training and post-training data.
All code and models of IXC-2.5-Reward and IXC-2.5-Chat are publicly available at https://github.com/InternLM/InternLM-XComposer.
Table 1: Overview of existing preference datasets used in IXC-2.5-Reward. IF denotes to Instruction Following.
Category Dataset
Text
IF General Tulu-3-IF-augmented-on-policy-8b [40]
UltraFeedback [17]
Safety hhh alignment [5], PKU-Safe [18]
SHP [24], Anthropic-hhrlhf [6]
Image
Chat WildVision-Battle [62]
General LLaVA-Critic [94], VL-Feedback [44],
RLAIF-V [101] MIA-DPO [54]
Table 2: Overview of the source of newly collected data used in IXC-2.5-Reward.
Category Dataset
Image
IF General in-house (will release)
KVQA [76], A-OKVQA [73], PMC-VQA [114]
Text-Rich AI2D [37], IconQA [56], TQA [38]
ChartQA [63], DVQA [36], ScienceQA [57]
Reasoning GeoQA [11], CLEVR-Math [47]
Super-CLEVR [45], TabMWP [58]
Video
General TrafficQA [96], FunQA [93], MiraData [35]
2 Related Work
Reward Model in Large Language Models
Reward models (RMs) are crucial for both Reinforcement Learning from Human Feedback (RLHF) [66, 7] and Test-time Scaling Laws [80, 29]. RMs have different implementation forms, such as (1) discriminative RM [8, 121, 48, 92, 88, 104, 55, 98], usually a sequence classifier that classifies input sequences into different categories, such as binary classification (“good” or “bad,”) or on a more granular scale [92, 88]. (2) generative RM [39, 105, 78, 91] that are prompted to generate the feedback in the form of text, often a critique or explanation of why a certain output is good or bad. (3) implicit RMs [32, 40] that are models optimized using DPO [70] that the predicted log probabilities are interpreted as implicit reward signal. Besides, RMs can also be divided into Outcome RMs (ORMs) [16] and Process RMs (PRMs) [86, 46, 75]. Our IXC-2.5-Reward belongs to the discriminative RM and ORM.
Reward Model in Large Vision-Language Models
Previous RMs for LVLMs [87, 94, 95] are limited to specific domains (e.g., reducing hallucination) or developed using relatively weak base models, which makes the implemented models significantly inferior to LLM RMs. The lack of effective multi-modal RMs has created a bottleneck in vision RLHF, forcing researchers to merely use the variants of the off-poly DPO algorithm [70]. Previous work using open-source LVLMs as generative RMs [102, 65, 95], injection of hallucinations with data augmentation techniques [22, 25, 120, 122, 67, 33, 21] and rule-based selection [9, 53] for DPO data selection, which potentially compromise performance compared to the on-policy RL solutions like PPO [71]. Moreover, lacking multi-modal RMs has also led to the reliance on human annotation [82, 100] or the use of proprietary models [109, 116] like GPT4 as generative RMs for DPO pair selection, which is expensive and unsustainable for large-scale applications. Although open-source RMs for LVLMs have lagged behind their LLM counterparts, the growing community interest highlights the need for multi-modal RMs, which motivates our work. In this work, we demonstrate that IXC-2.5-Reward is capable of combining with the PPO training and for DPO data selection at a low cost.
Reward Model Evaluations
The development of evaluation benchmarks is essential for improving RMs. Several comprehensive benchmarks have been proposed for evaluating RMs of LLMs, such as general abilities [41, 119, 51], multilingual [81, 28], RAG [34], and mathematical process reward [117]. The limited availability of multi-modal RMs has hampered the development of evaluation benchmarks, with existing benchmark [43] focusing solely on generative RMs and lacking the evaluation of process supervision. However, given the critical importance of RMs, we expect significant progress in multi-modal RM benchmarking in the future.
3 IXC2.5-Reward
Data Preparation
Reward models are trained using pairwise preference annotations (e.g., prompts $x$ with chosen responses $y_{c}$ and rejected responses $y_{r}$) that reflect human preferences. While existing public preference data is primarily textual, with limited image and scarce video examples, we train IXC-2.5-Reward using both open-source data and a newly collected dataset to ensure broader domain coverage.
Tab. 2 lists the open-source pairwise data used in IXC-2.5-Reward, primarily focused on instruction following, safety, and general knowledge. Tab. 2 details the source of our newly collected data, which is initially the supervised fine-tuning (SFT) data consisting of prompts $x$ and corresponding chosen responses $y_{c}$ across diverse domains: text-rich document understanding, math reasoning, and video understanding. We also collect some in-house data about the instruction following, which will be released in the future. To obtain rejected responses $y_{r}$, we prompt the SFT model, InternLM-XComposer-2.5 (IXC-2.5) [112] to generate multiple outputs for each prompt and then employ distinct selection criteria. For general and text-rich data, we use GPT-4o [31] with pairwise evaluation prompts to determine the rejected response that was evaluated worse than the SFT ground-truth answer. For math reasoning and instruction following data, we build verifier functions [40] that compare generated responses against ground-truth solutions to label the chosen and rejected data. Our newly collected data complements existing open-source data, creating a comprehensive, high-quality multi-modal preference dataset.
Model Architecture
Our reward model InternLM-XComposer 2.5-Reward (IXC-2.5-Reward) is built upon the SFT model (IXC-2.5) [111]. As shown in Fig. 1 (b), we use the pre-trained weights of IXC-2.5-Chat for most of the parts, such as the visual encoder and the MLP projector, which has aligned the image and video data with text modalities. Thus, the IXC-2.5-Reward is merely required to train preference data to predict the reward score and avoid using other pre-training data for modality alignment.
We replace the final linear layer of IXC-2.5 with a score head $f$ for IXC-2.5-Reward that predicts the reward score. Given the input prompt $x$ and the response $y$, the score head $f$ transforms the averaged hidden state features of all tokens into a binary scalar $r(x,y)$. This scalar value $r(x,y)$ serves as the predicted reward score for the inputs.
Loss Function
Our reward model is trained via the following loss function:
$\mathcal{L}_{\text{RM}}=-E(\log(\sigma(r(x,y_{w}))-r(x,y_{l}))),$ (1)
where $r(x,y_{w})$ and $r(x,y_{l})$ denotes to the reward score assigned to the prompt $x$ with the chosen data $y_{w}$ and rejected data $y_{l}$, respectively.
Training Strategy
As shown in Fig. 1 (b), we froze the model’s vision encoder and projector that are initialized from IXC-2.5 [112], training only the LLM (InternLM [112]) and the score head. Other components of IXC-2.5, such as the dynamic image partitioning mechanism for high-resolution inputs, remained unchanged.
Length Constraints
We remove data pairs where the length of the chosen response $y_{w}$ is significantly longer than the length of the rejected response $y_{l}$. This helps prevent the reward model from learning to associate length with quality. Notably, we found that the vulnerability of LLM-based evaluation to length bias, a known issue in LLMs [23], has also significant implications for LVLMs. Specifically, open-ended Visual Question Answering (VQA) benchmarks that employ LVLMs (e.g., GPT-4o) as judges are susceptible to inflated scores from overly long responses. Consequently, removing the length constraint on the reward model resulted in improved PPO policy performance. A detailed analysis is provided in Tab. 7.
Figure 2: Using IXC-2.5-Reward for Data Cleaning. We visualize the outlier and noisy examples detected by IXC-2.5-Reward with low reward scores from existing image and video instruction-tuning datasets, such as ALLaVA [10] and LLaVA-Video-178K [115]. The “Explain” refers to explanations of error causes as identified by human experts, rather than outputs generated by the reward model.
Table 3: Evaluation results on VLRewardBench [43]. The best and second-best results for proprietary models and open-source models are highlighted in bold and underlined, respectively.
Models #Param General Hallucination Reasoning Overall Acc Macro Acc
Proprietary Models
Gemini-1.5-Flash (2024-09-24) [83] - 47.8 59.6 58.4 57.6 55.3
Gemini-1.5-Pro (2024-09-24) [83] - 50.8 72.5 64.2 67.2 62.5
Claude-3.5-Sonnet (2024-06-22) [4] - 43.4 55.0 62.3 55.3 53.6
GPT-4o-mini (2024-07-18) [3] - 41.7 34.5 58.2 41.5 44.8
GPT-4o (2024-08-06) [3] - 49.1 67.6 70.5 65.8 62.4
Open-Source Models
LLaVA-OneVision-7B-ov [42] 7B 32.2 20.1 57.1 29.6 36.5
Qwen2-VL-7B [90] 7B 31.6 19.1 51.1 28.3 33.9
Molmo-7B [20] 7B 31.1 31.8 56.2 37.5 39.7
InternVL2-8B [85] 8B 35.6 41.1 59.0 44.5 45.2
LLaVA-Critic-8B [94] 8B 54.6 38.3 59.1 41.2 44.0
Llama-3.2-11B [84] 11B 33.3 38.4 56.6 42.9 42.8
Pixtral-12B [1] 12B 35.6 25.9 59.9 35.8 40.4
Molmo-72B [20] 72B 33.9 42.3 54.9 44.1 43.7
Qwen2-VL-72B [90] 72B 38.1 32.8 58.0 39.5 43.0
NVLM-D-72B [19] 72B 38.9 31.6 62.0 40.1 44.1
Llama-3.2-90B [84] 90B 42.6 57.3 61.7 56.2 53.9
IXC-2.5-Reward (Ours) 7B 84.7 62.5 62.9 65.8 70.0
4 The Applications of IXC-2.5-Reward
In this section, we further validate three applications of IXC-2.5-Reward for (1) RL training (Sec. 4.1), (2) test-time scaling (Sec. 4.2), and (3) data cleaning (Sec. 4.3).
4.1 IXC-2.5-Reward for RL training
Having the reward model IXC-2.5-Reward enables the application of on-policy reinforcement learning algorithms (e.g., PPO [71], RLOO [2], GRPO [77]) to optimize LVLM performance towards desired human preferences directly. Using the PPO [71] algorithm, we subsequently train the policy model (IXC-2.5-Chat, $\pi_{\theta}$) to maximize expected rewards from our reward model (IXC-2.5-Reward) while staying close to the reference model (IXC-2.5, $\pi_{\text{ref}}$) for stability. A critic model $V$, initialized from IXC-2.5-Reward, is trained alongside $\pi_{\theta}$ to reduce the variance of policy updates.
Data Prepration
Similar to findings in [30], we found that average reward scores differ across task domains (e.g., general, text-rich, reasoning). This work focuses on improving the policy model’s instruction following and open-ended chat abilities, which are critical for real-world applications such as stream chatting and human-AI interaction [110]. Simultaneously, we ensure that performance in other domains (e.g., text-rich, reasoning) is not degraded relative to the SFT model IXC-2.5. Using our multi-modal preference data (which trains IXC-2.5-Reward), we curate a prompt set that prioritizes general chat and instruction following, while ensuring diversity through the inclusion of text-rich documents, math reasoning, and video understanding.
PPO
The PPO training begins by sampling a prompt from our prompt set. Then, the policy $\theta_{\pi}$ model generates responses, and the reward model computes the reward score $r_{t}$ at each state $s_{t}$ at the time-step $t$. Given the reward score $r_{t}$ and and the critic model $V$, we compute the temporal difference error $\delta_{t}$, the Generalized Advantage Estimation (GAE) [72] $A_{t}$, and the Returns $R_{t}$ as:
$\delta_{t}$ $=r_{t}+\gamma\cdot V(s_{t+1})-V(s_{t}),$ (2)
$A_{t}$ $=\delta_{t}+\gamma\cdot\beta\cdot A_{t+1},$
$R_{t}$ $=A_{t}+V(s_{t}),$
where $\gamma$ is a discount factor that determines how much future rewards are valued compared to immediate rewards, and $\beta$ is the parameter that controls the trade-off between bias and variance in the advantage estimation. The advantage $A$ refers to how much better the policy model did than expected, and the returns $R$ is the cumulative reward.
Based on the advantage $A$, we compute the policy gradient loss $\mathcal{L}_{\text{PG}}$ to update the policy model $\pi_{\theta}$:
$\mathcal{L}_{\text{PG}}=\min(\frac{\pi_{\theta}}{\pi_{\text{ref}}}\cdot A,\text{clip}(\frac{\pi_{\theta}}{\pi_{\text{ref}}},1.0-\epsilon,1.0+\epsilon)\cdot A),$ (3)
where $\frac{\pi_{\theta}}{\pi_{\text{ref}}}$ is the log of the probability ratio between the policy model $\pi_{\theta}$ and the reference model $\pi_{\text{ref}}$, and $\epsilon$ is a hyper-parameter that controls the clipped ratio.
We further update the critic model via the Mean Squared Error (MSE) loss to minimize the difference between the predicted value of a state $V(s_{t})$ and the actual return $R_{t}$ obtained from state $t$:
$\mathcal{L}_{\text{critic}}=\sum_{t}\text{MSE}(V(s_{t}),R_{t}).$ (4)
In summary, with the help of IXC-2.5-Reward and PPO, we train the IXC-2.5-Chat to generate responses that improve the quality of multi-modal chat and follow user instructions. The quality of IXC-2.5-Chat also demonstrates the quality of IXC-2.5-Reward that provides the reward scores.
4.2 IXC-2.5-Reward for Test-Time Scaling
We further demonstrate that IXC-2.5-Reward is essential for scaling the inference-time capabilities of LVLMs. We select the Best-of-N (BoN) sampling technique that improves the quality of generated text by using the reward model. Specifically, the IXC-2.5-Chat model generates multiple ($N$) different text outputs with different random seeds for a given prompt. Subsequently, the reward model IXC-2.5-Reward scores each of these $N$ outputs, and the output with the highest score from the reward model is selected as the final output.
4.3 IXC-2.5-Reward for Data Cleaning
Garbage in, garbage out. Problematic samples in instruction tuning datasets negatively impact LVLM training. While existing methods [13] employ classifiers like CLIP [69] for filtering, these approaches have limitations, particularly with long-context inputs [108], high-resolution images, or videos. As shown in Fig. 2, we observe a strong correlation between low IXC-2.5-Reward scores and problematic samples, including hallucinations, empty answers, and irrelevant image/video-text pairings. Therefore, IXC-2.5-Reward effectively cleans both pre-training and post-training data for LVLMs.
5 Experiments
Table 4: Evaluation results on Reward Bench [41]. We report the performance of selective representative language-only RMs and previous multi-modal generative RMs.
Model Name LLM Chat Chat Hard Safety Reasoning Avg Score
Language-Only Reward Models
InternLM2-7B-Reward [8] InternLM2-7B 99.2 69.5 87.2 94.5 87.6
InternLM2-20B-Reward [8] InternLM2-20B 98.9 76.5 89.5 95.8 90.2
Skyword-Reward-Llama3.1-8B [48] Llama3.1-8B 95.8 87.3 90.8 96.2 92.5
INF-ORM-Llama3.1-70B [97] Llama3.1-70B 96.6 91.0 93.6 99.1 95.1
Multi-Modal Reward Models
QWen2-VL-7B [90] QWen2-7B 96.6 57.0 73.9 94.3 83.8
LLaVA-Critic-8B [94] LLaMA3-7B 96.9 52.8 81.7 83.5 80.0
IXC-2.5-Reward (Ours) InternL2-7B 90.8 83.8 87.8 90.0 88.6
Table 5: Evaluation results on RM-Bench [51]. We classify reward models into three types: sequence classifiers (Seq.), generative models, and implicit DPO models. Performance is reported across four domains (Chat, Math, Code, Safety) and three difficulty levels (Easy, Normal, Hard), along with average scores (Avg).
Model Name Type Chat Math Code Safety Easy Normal Hard Avg
Language-Only Reward Models
Tulu-2-dpo-13b [32] Implicit 66.4 51.4 51.8 85.4 86.9 66.7 37.7 63.8
InternLM2-7B-Reward [8] Seq. 61.7 71.4 49.7 85.5 85.4 70.7 45.1 67.1
InternLM2-20B-Reward [8] Seq. 63.1 66.8 56.7 86.5 82.6 71.6 50.7 68.3
Nemotron-4-340B-Reward [92] Generative 71.2 59.8 59.4 87.5 81.0 71.4 56.1 69.5
URM-LLaMa-3.1-8B [55] Seq. 71.2 61.8 54.1 93.1 84.0 73.2 53.0 70.0
Skyword-Reward-Llama3.1-8B [48] Seq. 69.5 60.6 54.5 95.7 89.0 74.7 46.6 70.1
Multi-Modal Reward Models
IXC-2.5-Reward (Ours) Seq. 65.5 55.9 51.7 93.8 87.5 71.3 47.4 68.8
Table 6: Evaluation results of our IXC-2.5-Chat model against previous SOTA proprietary and open-source models $\leq$10B (results are copied from [Uncaptioned image]OpenVLM Leaderboard and [Uncaptioned image]Open LMM Reasoning Leaderboard, accessed 01-Jan-2025). Best and second best results are highlighted.
Category Benchmark Evaluation Proprietary API Open-Source Model ($\leq$10B)
Previous-SOTA Previous-SOTA IXC-2.5 IXC-2.5-Chat
Instruction WildVision(0617) [62] Open 89.2 [31] 67.3 [94] 37.5 74.6
Following MIA${}_{(\text{val})}$ [68] Open 88.6 [31] 80.7 [90] 80.4 84.0
& Chat MM-MT${}_{(\text{val})}$ [1] Open 7.72 [31] 5.45 [90] 3.85 5.70
MM-Vet v2${}_{(\text{0613})}$ [103] Open 71.8 [4] 58.1 [14] 45.8 54.8
Knowledge MMBench${}_{(\text{v1.1})}$ [52] MCQ 85.7 [74] 82.7 [61] 79.4 79.0
MMMU${}_{(\text{val})}$ [106] MCQ 70.7 [31] 56.2 [14] 42.9 44.1
MMStar [12] MCQ 72.7 [74] 63.2 [14] 59.9 59.6
Reasoning MathVista${}_{(\text{mini})}$ [59] VQA 78.4 [74] 66.5 [60] 63.7 63.4
MathVerse${}_{(\text{vision-only})}$ [113] VQA 54.8 [26] 26.6 [50] 16.2 19.0
MathVision${}_{(\text{full})}$ [89] VQA 43.6 [26] 22.0 [50] 17.8 18.8
Text-Rich TextVQA${}_{(\text{val})}$ [79] VQA 82.0 [64] 78.5 [42] 78.2 81.3
ChartQA${}_{(\text{test})}$ [63] VQA 81.2 [64] 82.4 [99] 82.2 80.5
OCRBench [49] VQA 89.4 [74] 82.2 [14] 69.0 70.0
Figure 3: Visualizations of multi-modal dialogues generated by IXC-2.5-Chat on instruction following abilities.
Figure 4: Visualizations of multi-modal dialogues generated by IXC-2.5-Chat on open-ended questions.
Implementation Details
For IXC-2.5-Reward, the learning rates were set at 1e-5 with a batch size of 256. As for IXC-2.5-Chat, the learning rates were set at 5e-5 with a batch size of 256. We set the PPO hyper-parameters $\gamma=0.99$, $\beta=0.95$, and $\epsilon=0.2$.
5.1 Evaluation Results of IXC-2.5-Reward
Benchmarks
To evaluate IXC-2.5-Reward, we use diverse reward model benchmarks: (1) VL-RewardBench [43], encompassing 1250 multi-modal problems addressing general understanding, hallucination, and reasoning challenges; (2) Reward-Bench [41], with 2985 language-only problems including chat, chat hard, safety and reasoning; and (3) RM-Bench [51], comprising 1237 language-only problems across chat, math, code, and safety. RM-Bench defines three tracks (easy, normal, hard) that evaluate the sensitivity of reward models to subtle content variations and style biases. While Reward-Bench and RM-Bench are designed for reward models of language-only LLMs, we evaluate IXC-2.5-Reward on these benchmarks to demonstrate that our multi-modal reward model maintains strong language capabilities despite also processing image and video inputs.
5.1.1 Results on VL-RewardBench
Main Results
Tab. 3 presents the evaluation results of various multi-modal RMs on the VL-RewardBench [43]. Unlike previous multi-modal generative reward models, our IXC-2.5-Reward is a discriminative model that predicts a scalar reward. Our proposed IXC-2.5-Reward model, despite being an open-source 7B parameter model, outperforms all other open-source models. Notably, IXC-2.5-Reward achieves the highest overall accuracy (65.8%) among open-source models and the highest Macro Accuracy (70.0%) among all models, indicating its superior performance in handling diverse tasks within the VL-RewardBench.
Strong Performance on General Problems
The results in Table 3 reveal that IXC-2.5-Reward achieves a significantly higher accuracy (84.7%) on general problems compared to other generative RMs. We found the reason is attributed to these problems presenting a considerable challenge, often leading to tied judgments in previous LVLMs, whereas IXC-2.5-Reward demonstrates a greater ability to make correct classifications with different scalar scores.
5.1.2 Results on Reward Bench and RM-Bench
Main Results
We argue that multi-modal RMs should preserve strong language processing abilities despite the incorporation of image and video data during training. Consequently, we evaluate the performance of multi-modal reward models, including IXC-2.5-Reward, on Reward Bench (Tab. 4) and RM-Bench (Tab. 5). The results demonstrate that IXC-2.5-Reward achieves considerable performance and surpasses other multi-modal models on this benchmark.
Sensitivity to Content and Style
Consistent with findings in [51], IXC-2.5-Reward demonstrates sensitivity to subtle content variations and style biases, an issue often overlooked in multi-modal research. We believe further research is needed to enhance the robustness of multi-modal reward models.
5.2 Evaluation Results of IXC-2.5-Chat
Benchmarks
We select four representative benchmarks for evaluating the instruction following and in-the-wild chatting abilities of LVLMs. (1) The WildVision bench [62] uses prompts collected from user submissions, reflecting real-world multimodal interactions. (2) MIA-bench [68] that is specially designed to evaluate instruction following. (3) MM-MT [1] which is an instruction-following benchmark for multi-modal models, exhibits a strong correlation with LMSys-Vision ELO ratings [15]. (4) MM-Vet [103] that evaluate LVLMs on complex tasks such as language generation. These datasets contain open-ended questions and the referenced answers, and evaluation is performed using an LLM-as-a-Judge [118] approach, which involves using a judge model like GPT-4o [31] to predict scores.
We also report the performance on other categories, such as MMBench [52], MMMU [106] and MMStar [12] for general knowledge, MathVerse [113] and MathVision [89] for math reasoning, TextVQA [79], ChartQA [63] and OCRbench [49] for text-rich document understanding. These benchmarks utilize multiple-choice questions (MCQ) or visual question answering (VQA), where responses are limited to short keywords and evaluated based on string matching.
Results on Instruction Following & Chat
Tab. 6 shows that IXC-2.5-Chat outperforms previous SOTA models across multiple benchmarks (WildVision, MIA, and MM-MT), demonstrating significant improvements in multi-modal understanding with instruction following ability and provide more comprehensive information for in-the-wild chat scenario.
Results on Other Categories
On other categories (Knowledge, Reasoning, and Text-Rich), IXC-2.5-Chat performs comparably to the supervised fine-tuned (SFT) model IXC-2.5, demonstrating that RL training with IXC-2.5-Reward improves instruction following and conversational ability without sacrificing performance in these areas.
Table 7: Ablation Studies of the impact of response length constraints of reward models that guided training IXC-2.5-Chat.
Avg Wild MIA MM-MT MM-Vet
Tokens Vision v2
w/o Length Constraints 361 76.2 87.0 5.86 56.6
IXC-2.5-Chat 274 74.6 84.0 5.70 54.8
The Impact of Length Constraints
To prevent the chat model from generating overly long responses to artificially inflate rewards, we introduce length constraints on the ratio of chosen to rejected responses during training reward model IXC-2.5-Reward. The ablation study results of length constraints are present in Tab. 7. On the WildVision benchmark, we compute the average token length of the model’s responses. We observe a substantial increase in average token length, from 274 to 361, when length constraints were not applied. Surprisingly, removing length constraints yielded substantial improvements in open-ended benchmarks, achieving state-of-the-art results. Such observation is because these benchmarks do not penalize length in their evaluation prompts, judge models (e.g., GPT-4) tend to favor longer responses, even if they contain unnecessary details that detract from the user experience. As our focus is on optimizing user experience, not benchmark scores, we retain the length constraints. Following the precedent set by language-only benchmarks (e.g., [23]), we believe multi-modal Chat Arena and dialogue benchmarks should also address potential length and style biases in their evaluation protocols in future work.
Results on Test-Time Scaling
According to Tab. 8, we observe that the Best-of-$N$ sampling further improves the results. The averaged tokens is increased slightly (from 274 to 283), demonstrate that the improvements is bring from the high-quality response, rather than hacking the length bias in Tab. 7.
Table 8: Results of Best-of-$N$ (BoN) sampling for test-time scaling with IXC-2.5-Reward.
N Avg Wild MIA MM-MT MM-Vet
Tokens Vision v2
IXC-2.5-Chat 274 74.6 84.0 5.70 54.8
IXC-2.5-Chat + BoN 4 283 77.7 87.3 6.03 56.3
Visualization Results
We present the visualization examples of IXC-2.5-Chat on a series of topics, such as instruction following (Fig. 3) and open-ended questions (Fig. 4). These figures reveal that IXC-2.5-Chat demonstrates several key advantages, including superior organization and presentation, more comprehensive and in-depth answers, and more detailed explanations. These strengths significantly enhance IXC-2.5-Chat’s effectiveness in multi-modal chat interactions.
6 Conclusion and Future Work
We present IXC-2.5-Reward, a multi-modal reward model that is capable of multi-modal RL training, test-time scaling and data cleaning. IXC-2.5-Reward supports text, image, and video inputs and is trained on diverse, high-quality preference data to ensure robust preference discrimination. Using IXC-2.5-Reward, we further trained IXC-2.5-Chat via RLHF techniques to optimize the multi-modal user chat experience, focusing on providing detailed explanations and in-depth answers.
We believe that exploring multi-modal reward models with on-policy reinforcement learning algorithms holds significant promise for future research. For instance, while IXC-2.5-Reward can score multi-image and video inputs, existing reward benchmarks and applying RL algorithms for video alignment remain scarce. Moreover, extending the development of multi-modal reward models beyond chat to other domains, such as process reward models (PRMs) for multi-modal reasoning, is a promising direction.

### v2

@@
 We integrate IXC-2.5-Reward with Proximal Policy Optimization (PPO) yields IXC-2.5-Chat, which shows consistent improvements in instruction following and multi-modal open-ended dialogue; (2) Selecting the best response from candidate responses for test-time scaling; and (3) Filtering outlier or noisy samples from existing image and video instruction tuning training data.
-To ensure reproducibility and facilitate further research, we have open-sourced all model weights and training recipes at https://github.com/InternLM/InternLM-XComposer.
+InternLM-XComposer2.5-Reward: A Simple Yet Effective
+Multi-Modal Reward Model
 […]
-1 Introduction
-If you don’t know where you are going, you’ll end up some
-place else.
-— Yogi Berra
+[…]
 Figure 1: (a) To train the IXC-2.5-Reward, we construct a multi-modal preference dataset spanning diverse domains (e.g., natural scenes, text-rich, reasoning) and modalities (image, text, video).
@@
 (c) The IXC-2.5-Reward guides policy training for IXC-2.5-Chat via reinforcement learning.
-Reward Models (RMs) [8, 121, 48, 92, 88, 104, 55, 98, 105, 78, 91] provide the crucial direction guidance about how well an AI model’s outputs align with human preference, and benefit Large Language Models (LLMs) in training and inference.
-During training, RMs are often used with reinforcement learning from human feedback (RLHF) [66, 7, 71, 70] to penalize undesirable model behaviors and encourage outputs that align with human values.
-At inference, RMs facilitate test-time scaling strategies [80, 27], such as selecting the best response from candidate outputs or providing step-by-step critiques for complex reasoning tasks [107, 29].
+1 Introduction
+If you don’t know where you are going, you’ll end
+up some place else.
+— Yogi Berra
+Reward Models (RMs) Cai et al.
+(2024); Zhu et al.
+(2023); Liu et al.
+(2024a); Wang et al.
+(2024f, b); Yuan et al.
+(2024a); Lou et al.
+(2024); Yang et al.
+(2024b); Yuan et al.
+(2024b); Shiwen et al.
+(2024); Wang et al.
+(2024e) provide the crucial direction guidance about how well an AI model’s outputs align with human preference, and benefit Large Language Models (LLMs) in training and inference.
+During training, RMs are often used with reinforcement learning from human feedback (RLHF) Ouyang et al.
+(2022); Bai et al.
+(2022b); Schulman et al.
+(2017); Rafailov et al.
+(2024) to penalize undesirable model behaviors and encourage outputs that align with human values.
+At inference, RMs facilitate test-time scaling strategies Snell et al.
+(2024); Gulcehre et al.
+(2023), such as selecting the best response from candidate outputs or providing step-by-step critiques for complex reasoning tasks Zelikman et al.
+(2022); Hosseini et al.
+(2024).
 Despite their crucial role in both training and inference, multi-modal RMs for Large Vision Language Models (LVLMs) remain notably underexplored compared to language-only RMs for LLMs.
 Because current preference data is predominantly text-based and skewed toward specific domains (e.g., safety), data scarcity poses a significant challenge to training multi-modal RMs for diverse modalities such as images, videos, and text.
-Consequently, existing multi-modal RMs [87, 95] are largely constrained to narrow domains (e.g., mitigating hallucinations) or rely on prompting LVLMs with evaluation prompts, effectively functioning as generative RMs [94].
+Consequently, existing multi-modal RMs Wang et al.
+(2024a); Xiyao et al.
+(2024) are largely constrained to narrow domains (e.g., mitigating hallucinations) or rely on prompting LVLMs with evaluation prompts, effectively functioning as generative RMs Xiong et al.
+(2024).
 The limitation of multi-modal RMs subsequently constrains the capabilities of open-source LVLMs such as instruction following and safety-should-refuse, thereby hampering user interaction experience in multi-modal chat scenarios.
@@
 1(a)) to construct multi-modal preference data, and also incorporate existing high-quality datasets.
-This pipeline selects prompts across diverse domains for text, image, and video inputs, generates corresponding responses, and then uses GPT-4o [31] or verifier [40] to perform preference judgments.
+This pipeline selects prompts across diverse domains for text, image, and video inputs, generates corresponding responses, and then uses GPT-4o Hurst et al.
+(2024) or verifier Lambert et al.
+(2024a) to perform preference judgments.
 Trained on our preference data, IXC-2.5-Reward effectively evaluates both visual (image and video) and textual inputs (Fig.
 1 (b)).
-IXC-2.5-Reward achieves best performance on multi-modal VL-RewardBench [43] (70.0%) that beat all previous generative RMs including Gemini-1.5-Pro (62.5%) and GPT-4o (62.4%).
-Even on uni-modal (text) RM benchmarks, IXC-2.5-Reward also demonstrates good results, with an average score of $88.6\%$ on Reward-Bench [41] and $68.8\%$ on RM-Bench [51].
+IXC-2.5-Reward achieves best performance on multi-modal VL-RewardBench Li et al.
+(2024b) (70.0%) that beat all previous generative RMs including Gemini-1.5-Pro (62.5%) and GPT-4o (62.4%).
+Even on uni-modal (text) RM benchmarks, IXC-2.5-Reward also demonstrates good results, with an average score of $88.6\%$ on Reward-Bench Lambert et al.
+(2024b) and $68.8\%$ on RM-Bench Liu et al.
+(2024b).
 We further demonstrate the effectiveness of IXC-2.5-Reward in the following three aspects:
@@
 This suggests that IXC-2.5-Reward can effectively clean LVLM pre-training and post-training data.
-All code and models of IXC-2.5-Reward and IXC-2.5-Chat are publicly available at https://github.com/InternLM/InternLM-XComposer.
-Table 1: Overview of existing preference datasets used in IXC-2.5-Reward. IF denotes to Instruction Following.
+Table 1: Overview of existing preference datasets used in IXC-2.5-Reward.
 Category Dataset
 Text
-IF General Tulu-3-IF-augmented-on-policy-8b [40]
-UltraFeedback [17]
-Safety hhh alignment [5], PKU-Safe [18]
-SHP [24], Anthropic-hhrlhf [6]
+IF General Tulu-3-IF-augmented-on-policy-8b Lambert et al. (2024a)
+UltraFeedback Cui et al. (2024)
+Safety hhh alignment Askell et al. (2021), PKU-Safe Dai et al. (2024a)
+SHP Ethayarajh et al. (2022), Anthropic-hhrlhf Bai et al. (2022a)
 Image
-Chat WildVision-Battle [62]
-General LLaVA-Critic [94], VL-Feedback [44],
-RLAIF-V [101] MIA-DPO [54]
+Chat WildVision-Battle Lu et al. (2024c)
+General LLaVA-Critic Xiong et al. (2024), VL-Feedback Li et al. (2024c),
+RLAIF-V Yu et al. (2024b) MIA-DPO Liu et al. (2024e)
 Table 2: Overview of the source of newly collected data used in IXC-2.5-Reward.
@@
 Image
-IF General in-house (will release)
-KVQA [76], A-OKVQA [73], PMC-VQA [114]
-Text-Rich AI2D [37], IconQA [56], TQA [38]
-ChartQA [63], DVQA [36], ScienceQA [57]
-Reasoning GeoQA [11], CLEVR-Math [47]
-Super-CLEVR [45], TabMWP [58]
+IF General MM-IFDPO-23k Ding et al. (2025)
+KVQA Shah et al. (2019), A-OKVQA Schwenk et al. (2022), PMC-VQA Zhang et al. (2023)
+Text-Rich AI2D Kembhavi et al. (2016), IconQA Lu et al. (2021), TQA Kembhavi et al. (2017)
+ChartQA Masry et al. (2022), DVQA Kafle et al. (2018), ScienceQA Lu et al. (2022a)
+Reasoning GeoQA Chen et al. (2021), CLEVR-Math Lindström and Abraham (2022)
+Super-CLEVR Li et al. (2023), TabMWP Lu et al. (2022b)
 Video
-General TrafficQA [96], FunQA [93], MiraData [35]
+General TrafficQA Xu et al. (2021), FunQA Xie et al. (2024), MiraData Ju et al. (2024)
 2 Related Work
-Reward Model in Large Language Models
-Reward models (RMs) are crucial for both Reinforcement Learning from Human Feedback (RLHF) [66, 7] and Test-time Scaling Laws [80, 29].
-RMs have different implementation forms, such as (1) discriminative RM [8, 121, 48, 92, 88, 104, 55, 98], usually a sequence classifier that classifies input sequences into different categories, such as binary classification (“good” or “bad,”) or on a more granular scale [92, 88].
-(2) generative RM [39, 105, 78, 91] that are prompted to generate the feedback in the form of text, often a critique or explanation of why a certain output is good or bad.
-(3) implicit RMs [32, 40] that are models optimized using DPO [70] that the predicted log probabilities are interpreted as implicit reward signal.
-Besides, RMs can also be divided into Outcome RMs (ORMs) [16] and Process RMs (PRMs) [86, 46, 75].
+Reward Model in Large Language Models.
+Reward models (RMs) are crucial for both Reinforcement Learning from Human Feedback (RLHF) Ouyang et al.
+(2022); Bai et al.
+(2022b) and Test-time Scaling Laws Snell et al.
+(2024); Hosseini et al.
+(2024).
+RMs have different implementation forms, such as (1) discriminative RM Cai et al.
+(2024); Zhu et al.
+(2023); Liu et al.
+(2024a); Wang et al.
+(2024f, b); Yuan et al.
+(2024a); Lou et al.
+(2024); Yang et al.
+(2024b), usually a sequence classifier that classifies input sequences into different categories, such as binary classification (“good” or “bad,”) or on a more granular scale Wang et al.
+(2024f, b).
+(2) generative RM Kim et al.
+(2023); Yuan et al.
+(2024b); Shiwen et al.
+(2024); Wang et al.
+(2024e) that are prompted to generate the feedback in the form of text, often a critique or explanation of why a certain output is good or bad.
+(3) implicit RMs Ivison et al.
+(2023); Lambert et al.
+(2024a) that are models optimized using DPO Rafailov et al.
+(2024) that the predicted log probabilities are interpreted as implicit reward signal.
+Besides, RMs can also be divided into Outcome RMs (ORMs) Cobbe et al.
+(2021) and Process RMs (PRMs) Uesato et al.
+(2022); Lightman et al.
+(2023); Setlur et al.
+(2024).
 Our IXC-2.5-Reward belongs to the discriminative RM and ORM.
-Reward Model in Large Vision-Language Models
-Previous RMs for LVLMs [87, 94, 95] are limited to specific domains (e.g., reducing hallucination) or developed using relatively weak base models, which makes the implemented models significantly inferior to LLM RMs.
-The lack of effective multi-modal RMs has created a bottleneck in vision RLHF, forcing researchers to merely use the variants of the off-poly DPO algorithm [70].
-Previous work using open-source LVLMs as generative RMs [102, 65, 95], injection of hallucinations with data augmentation techniques [22, 25, 120, 122, 67, 33, 21] and rule-based selection [9, 53] for DPO data selection, which potentially compromise performance compared to the on-policy RL solutions like PPO [71].
-Moreover, lacking multi-modal RMs has also led to the reliance on human annotation [82, 100] or the use of proprietary models [109, 116] like GPT4 as generative RMs for DPO pair selection, which is expensive and unsustainable for large-scale applications.
+Reward Model in Large Vision-Language Models.
+Previous RMs for LVLMs Wang et al.
+(2024a); Xiong et al.
+(2024); Xiyao et al.
+(2024) are limited to specific domains (e.g., reducing hallucination) or developed using relatively weak base models, which makes the implemented models significantly inferior to LLM RMs.
+The lack of effective multi-modal RMs has created a bottleneck in vision RLHF, forcing researchers to merely use the variants of the off-poly DPO algorithm Rafailov et al.
+(2024).
+Previous work using open-source LVLMs as generative RMs Yu et al.
+(2024c); Ouali et al.
+(2025); Xiyao et al.
+(2024), injection of hallucinations with data augmentation techniques Deng et al.
+(2024b); Favero et al.
+(2024); Zhou et al.
+(2024b); Zhu et al.
+(2024); Pi et al.
+(2025); Jiang et al.
+(2024); Deng et al.
+(2024a) and rule-based selection Cao et al.
+(2024); Liu et al.
+(2024f) for DPO data selection, which potentially compromise performance compared to the on-policy RL solutions like PPO Schulman et al.
+(2017).
+Moreover, lacking multi-modal RMs has also led to the reliance on human annotation Sun et al.
+(2023); Yu et al.
+(2024a) or the use of proprietary models Zhang et al.
+(2024a); Zhao et al.
+(2023) like GPT4 as generative RMs for DPO pair selection, which is expensive and unsustainable for large-scale applications.
 Although open-source RMs for LVLMs have lagged behind their LLM counterparts, the growing community interest highlights the need for multi-modal RMs, which motivates our work.
 In this work, we demonstrate that IXC-2.5-Reward is capable of combining with the PPO training and for DPO data selection at a low cost.
-Reward Model Evaluations
+Reward Model Evaluations.
 The development of evaluation benchmarks is essential for improving RMs.
-Several comprehensive benchmarks have been proposed for evaluating RMs of LLMs, such as general abilities [41, 119, 51], multilingual [81, 28], RAG [34], and mathematical process reward [117].
-The limited availability of multi-modal RMs has hampered the development of evaluation benchmarks, with existing benchmark [43] focusing solely on generative RMs and lacking the evaluation of process supervision.
+Several comprehensive benchmarks have been proposed for evaluating RMs of LLMs, such as general abilities Lambert et al.
+(2024b); Zhou et al.
+(2024a); Liu et al.
+(2024b), multilingual Son et al.
+(2024); Gureja et al.
+(2024), RAG Jin et al.
+(2024), and mathematical process reward Zheng et al.
+(2024).
+The limited availability of multi-modal RMs has hampered the development of evaluation benchmarks, with existing benchmark Li et al.
+(2024b) focusing solely on generative RMs and lacking the evaluation of process supervision.
 However, given the critical importance of RMs, we expect significant progress in multi-modal RM benchmarking in the future.
 3 IXC2.5-Reward
-Data Preparation
+Data Preparation.
 Reward models are trained using pairwise preference annotations (e.g., prompts $x$ with chosen responses $y_{c}$ and rejected responses $y_{r}$) that reflect human preferences.
@@
 We also collect some in-house data about the instruction following, which will be released in the future.
-To obtain rejected responses $y_{r}$, we prompt the SFT model, InternLM-XComposer-2.5 (IXC-2.5) [112] to generate multiple outputs for each prompt and then employ distinct selection criteria.
-For general and text-rich data, we use GPT-4o [31] with pairwise evaluation prompts to determine the rejected response that was evaluated worse than the SFT ground-truth answer.
-For math reasoning and instruction following data, we build verifier functions [40] that compare generated responses against ground-truth solutions to label the chosen and rejected data.
+To obtain rejected responses $y_{r}$, we prompt the SFT model, InternLM-XComposer-2.5 (IXC-2.5) Zhang et al.
+(2024c) to generate multiple outputs for each prompt and then employ distinct selection criteria.
+For general and text-rich data, we use GPT-4o Hurst et al.
+(2024) with pairwise evaluation prompts to determine the rejected response that was evaluated worse than the SFT ground-truth answer.
+For math reasoning and instruction following data, we build verifier functions Lambert et al.
+(2024a) that compare generated responses against ground-truth solutions to label the chosen and rejected data.
 Our newly collected data complements existing open-source data, creating a comprehensive, high-quality multi-modal preference dataset.
-Model Architecture
-Our reward model InternLM-XComposer 2.5-Reward (IXC-2.5-Reward) is built upon the SFT model (IXC-2.5) [111].
+Model Architecture.
+Our reward model InternLM-XComposer 2.5-Reward (IXC-2.5-Reward) is built upon the SFT model (IXC-2.5) Zhang et al.
+(2024d).
 As shown in Fig.
@@
 This scalar value $r(x,y)$ serves as the predicted reward score for the inputs.
-Loss Function
-Our reward model is trained via the following loss function:
-$\mathcal{L}_{\text{RM}}=-E(\log(\sigma(r(x,y_{w}))-r(x,y_{l}))),$ (1)
+Loss Function. Our reward model is trained via the following loss function:
+$\mathcal{L}_{\text{RM}}=-E(\log(\sigma(r(x,y_{w})-r(x,y_{l})))),$ (1)
 where $r(x,y_{w})$ and $r(x,y_{l})$ denotes to the reward score assigned to the prompt $x$ with the chosen data $y_{w}$ and rejected data $y_{l}$, respectively.
-Training Strategy
+Training Strategy.
 As shown in Fig.
-1 (b), we froze the model’s vision encoder and projector that are initialized from IXC-2.5 [112], training only the LLM (InternLM [112]) and the score head.
+1 (b), we froze the model’s vision encoder and projector that are initialized from IXC-2.5 Zhang et al.
+(2024c), training only the LLM (InternLM Zhang et al.
+(2024c)) and the score head.
 Other components of IXC-2.5, such as the dynamic image partitioning mechanism for high-resolution inputs, remained unchanged.
-Length Constraints
+Length Constraints.
 We remove data pairs where the length of the chosen response $y_{w}$ is significantly longer than the length of the rejected response $y_{l}$.
 This helps prevent the reward model from learning to associate length with quality.
-Notably, we found that the vulnerability of LLM-based evaluation to length bias, a known issue in LLMs [23], has also significant implications for LVLMs.
+Notably, we found that the vulnerability of LLM-based evaluation to length bias, a known issue in LLMs Dubois et al.
+(2024), has also significant implications for LVLMs.
 Specifically, open-ended Visual Question Answering (VQA) benchmarks that employ LVLMs (e.g., GPT-4o) as judges are susceptible to inflated scores from overly long responses.
@@
 Figure 2: Using IXC-2.5-Reward for Data Cleaning.
-We visualize the outlier and noisy examples detected by IXC-2.5-Reward with low reward scores from existing image and video instruction-tuning datasets, such as ALLaVA [10] and LLaVA-Video-178K [115].
+We visualize the outlier and noisy examples detected by IXC-2.5-Reward with low reward scores from existing image and video instruction-tuning datasets, such as ALLaVA Chen et al.
+(2024a) and LLaVA-Video-178K Zhang et al.
+(2024e).
 The “Explain” refers to explanations of error causes as identified by human experts, rather than outputs generated by the reward model.
-Table 3: Evaluation results on VLRewardBench [43]. The best and second-best results for proprietary models and open-source models are highlighted in bold and underlined, respectively.
+Table 3: Evaluation results on VLRewardBench Li et al. (2024b). The best and second-best results for proprietary models and open-source models are highlighted in bold and underlined, respectively.
 Models #Param General Hallucination Reasoning Overall Acc Macro Acc
 Proprietary Models
-Gemini-1.5-Flash (2024-09-24) [83] - 47.8 59.6 58.4 57.6 55.3
-Gemini-1.5-Pro (2024-09-24) [83] - 50.8 72.5 64.2 67.2 62.5
-Claude-3.5-Sonnet (2024-06-22) [4] - 43.4 55.0 62.3 55.3 53.6
-GPT-4o-mini (2024-07-18) [3] - 41.7 34.5 58.2 41.5 44.8
-GPT-4o (2024-08-06) [3] - 49.1 67.6 70.5 65.8 62.4
+Gemini-1.5-Flash (2024-09-24) Team (2024a) - 47.8 59.6 58.4 57.6 55.3
+Gemini-1.5-Pro (2024-09-24) Team (2024a) - 50.8 72.5 64.2 67.2 62.5
+Claude-3.5-Sonnet (2024-06-22) Anthropic (2024) - 43.4 55.0 62.3 55.3 53.6
+GPT-4o-mini (2024-07-18) AI (2024) - 41.7 34.5 58.2 41.5 44.8
+GPT-4o (2024-08-06) AI (2024) - 49.1 67.6 70.5 65.8 62.4
 Open-Source Models
-LLaVA-OneVision-7B-ov [42] 7B 32.2 20.1 57.1 29.6 36.5
-Qwen2-VL-7B [90] 7B 31.6 19.1 51.1 28.3 33.9
-Molmo-7B [20] 7B 31.1 31.8 56.2 37.5 39.7
-InternVL2-8B [85] 8B 35.6 41.1 59.0 44.5 45.2
-LLaVA-Critic-8B [94] 8B 54.6 38.3 59.1 41.2 44.0
-Llama-3.2-11B [84] 11B 33.3 38.4 56.6 42.9 42.8
-Pixtral-12B [1] 12B 35.6 25.9 59.9 35.8 40.4
-Molmo-72B [20] 72B 33.9 42.3 54.9 44.1 43.7
-Qwen2-VL-72B [90] 72B 38.1 32.8 58.0 39.5 43.0
-NVLM-D-72B [19] 72B 38.9 31.6 62.0 40.1 44.1
-Llama-3.2-90B [84] 90B 42.6 57.3 61.7 56.2 53.9
+LLaVA-OneVision-7B-ov Li et al. (2024a) 7B 32.2 20.1 57.1 29.6 36.5
+Qwen2-VL-7B Wang et al. (2024d) 7B 31.6 19.1 51.1 28.3 33.9
+Molmo-7B Deitke et al. (2024) 7B 31.1 31.8 56.2 37.5 39.7
+InternVL2-8B Team (2024c) 8B 35.6 41.1 59.0 44.5 45.2
+LLaVA-Critic-8B Xiong et al. (2024) 8B 54.6 38.3 59.1 41.2 44.0
+Llama-3.2-11B Team (2024b) 11B 33.3 38.4 56.6 42.9 42.8
+Pixtral-12B Agrawal et al. (2024) 12B 35.6 25.9 59.9 35.8 40.4
+Molmo-72B Deitke et al. (2024) 72B 33.9 42.3 54.9 44.1 43.7
+Qwen2-VL-72B Wang et al. (2024d) 72B 38.1 32.8 58.0 39.5 43.0
+NVLM-D-72B Dai et al. (2024b) 72B 38.9 31.6 62.0 40.1 44.1
+Llama-3.2-90B Team (2024b) 90B 42.6 57.3 61.7 56.2 53.9
 IXC-2.5-Reward (Ours) 7B 84.7 62.5 62.9 65.8 70.0
@@
 4.1 IXC-2.5-Reward for RL training
-Having the reward model IXC-2.5-Reward enables the application of on-policy reinforcement learning algorithms (e.g., PPO [71], RLOO [2], GRPO [77]) to optimize LVLM performance towards desired human preferences directly.
-Using the PPO [71] algorithm, we subsequently train the policy model (IXC-2.5-Chat, $\pi_{\theta}$) to maximize expected rewards from our reward model (IXC-2.5-Reward) while staying close to the reference model (IXC-2.5, $\pi_{\text{ref}}$) for stability.
+Having the reward model IXC-2.5-Reward enables the application of on-policy reinforcement learning algorithms (e.g., PPO Schulman et al.
+(2017), RLOO Ahmadian et al.
+(2024), GRPO Shao et al.
+(2024)) to optimize LVLM performance towards desired human preferences directly.
+Using the PPO Schulman et al.
+(2017) algorithm, we subsequently train the policy model (IXC-2.5-Chat, $\pi_{\theta}$) to maximize expected rewards from our reward model (IXC-2.5-Reward) while staying close to the reference model (IXC-2.5, $\pi_{\text{ref}}$) for stability.
 A critic model $V$, initialized from IXC-2.5-Reward, is trained alongside $\pi_{\theta}$ to reduce the variance of policy updates.
-Data Prepration
-Similar to findings in [30], we found that average reward scores differ across task domains (e.g., general, text-rich, reasoning).
-This work focuses on improving the policy model’s instruction following and open-ended chat abilities, which are critical for real-world applications such as stream chatting and human-AI interaction [110].
+Data Prepration.
+Similar to findings in Hou et al.
+(2024), we found that average reward scores differ across task domains (e.g., general, text-rich, reasoning).
+This work focuses on improving the policy model’s instruction following and open-ended chat abilities, which are critical for real-world applications such as stream chatting and human-AI interaction Zhang et al.
+(2024b).
 Simultaneously, we ensure that performance in other domains (e.g., text-rich, reasoning) is not degraded relative to the SFT model IXC-2.5.
 Using our multi-modal preference data (which trains IXC-2.5-Reward), we curate a prompt set that prioritizes general chat and instruction following, while ensuring diversity through the inclusion of text-rich documents, math reasoning, and video understanding.
-PPO
+PPO.
 The PPO training begins by sampling a prompt from our prompt set.
 Then, the policy $\theta_{\pi}$ model generates responses, and the reward model computes the reward score $r_{t}$ at each state $s_{t}$ at the time-step $t$.
-Given the reward score $r_{t}$ and and the critic model $V$, we compute the temporal difference error $\delta_{t}$, the Generalized Advantage Estimation (GAE) [72] $A_{t}$, and the Returns $R_{t}$ as:
+Given the reward score $r_{t}$ and and the critic model $V$, we compute the temporal difference error $\delta_{t}$, the Generalized Advantage Estimation (GAE) Schulman et al.
+(2018) $A_{t}$, and the Returns $R_{t}$ as:
 $\delta_{t}$ $=r_{t}+\gamma\cdot V(s_{t+1})-V(s_{t}),$ (2)
@@
 Problematic samples in instruction tuning datasets negatively impact LVLM training.
-While existing methods [13] employ classifiers like CLIP [69] for filtering, these approaches have limitations, particularly with long-context inputs [108], high-resolution images, or videos.
+While existing methods Chen et al.
+(2024c) employ classifiers like CLIP Radford et al.
+(2021) for filtering, these approaches have limitations, particularly with long-context inputs Zhang et al.
+(2025a), high-resolution images, or videos.
 As shown in Fig.
@@
 5 Experiments
-Table 4: Evaluation results on Reward Bench [41]. We report the performance of selective representative language-only RMs and previous multi-modal generative RMs.
+Table 4: Evaluation results on Reward Bench Lambert et al. (2024b). We report the performance of selective representative language-only RMs and previous multi-modal generative RMs.
 Model Name LLM Chat Chat Hard Safety Reasoning Avg Score
 Language-Only Reward Models
-InternLM2-7B-Reward [8] InternLM2-7B 99.2 69.5 87.2 94.5 87.6
-InternLM2-20B-Reward [8] InternLM2-20B 98.9 76.5 89.5 95.8 90.2
-Skyword-Reward-Llama3.1-8B [48] Llama3.1-8B 95.8 87.3 90.8 96.2 92.5
-INF-ORM-Llama3.1-70B [97] Llama3.1-70B 96.6 91.0 93.6 99.1 95.1
+InternLM2-7B-Reward Cai et al. (2024) InternLM2-7B 99.2 69.5 87.2 94.5 87.6
+InternLM2-20B-Reward Cai et al. (2024) InternLM2-20B 98.9 76.5 89.5 95.8 90.2
+Skyword-Reward-Llama3.1-8B Liu et al. (2024a) Llama3.1-8B 95.8 87.3 90.8 96.2 92.5
+INF-ORM-Llama3.1-70B Yang et al. (2024a) Llama3.1-70B 96.6 91.0 93.6 99.1 95.1
 Multi-Modal Reward Models
-QWen2-VL-7B [90] QWen2-7B 96.6 57.0 73.9 94.3 83.8
-LLaVA-Critic-8B [94] LLaMA3-7B 96.9 52.8 81.7 83.5 80.0
-IXC-2.5-Reward (Ours) InternL2-7B 90.8 83.8 87.8 90.0 88.6
-Table 5: Evaluation results on RM-Bench [51].
+QWen2-VL-7B Wang et al. (2024d) QWen2-7B 96.6 57.0 73.9 94.3 83.8
+LLaVA-Critic-8B Xiong et al. (2024) LLaMA3-7B 96.9 52.8 81.7 83.5 80.0
+IXC-2.5-Reward (Ours) InternLM2-7B 90.8 83.8 87.8 90.0 88.6
+Table 5: Evaluation results on RM-Bench Liu et al.
+(2024b).
 We classify reward models into three types: sequence classifiers (Seq.), generative models, and implicit DPO models.
-Performance is reported across four domains (Chat, Math, Code, Safety) and three difficulty levels (Easy, Normal, Hard), along with average scores (Avg).
+Performance is reported across four domains (Chat, Math, Code, Safety) and three difficulty levels (Easy, Normal, Hard), along with average scores.
 Model Name Type Chat Math Code Safety Easy Normal Hard Avg
 Language-Only Reward Models
-Tulu-2-dpo-13b [32] Implicit 66.4 51.4 51.8 85.4 86.9 66.7 37.7 63.8
-InternLM2-7B-Reward [8] Seq. 61.7 71.4 49.7 85.5 85.4 70.7 45.1 67.1
-InternLM2-20B-Reward [8] Seq. 63.1 66.8 56.7 86.5 82.6 71.6 50.7 68.3
-Nemotron-4-340B-Reward [92] Generative 71.2 59.8 59.4 87.5 81.0 71.4 56.1 69.5
-URM-LLaMa-3.1-8B [55] Seq. 71.2 61.8 54.1 93.1 84.0 73.2 53.0 70.0
-Skyword-Reward-Llama3.1-8B [48] Seq. 69.5 60.6 54.5 95.7 89.0 74.7 46.6 70.1
+Tulu-2-dpo-13b Ivison et al. (2023) Implicit 66.4 51.4 51.8 85.4 86.9 66.7 37.7 63.8
+InternLM2-7B-Reward Cai et al. (2024) Seq. 61.7 71.4 49.7 85.5 85.4 70.7 45.1 67.1
+InternLM2-20B-Reward Cai et al. (2024) Seq. 63.1 66.8 56.7 86.5 82.6 71.6 50.7 68.3
+Nemotron-4-340B-Reward Wang et al. (2024f) Generative 71.2 59.8 59.4 87.5 81.0 71.4 56.1 69.5
+URM-LLaMa-3.1-8B Lou et al. (2024) Seq. 71.2 61.8 54.1 93.1 84.0 73.2 53.0 70.0
+Skyword-Reward-Llama3.1-8B Liu et al. (2024a) Seq. 69.5 60.6 54.5 95.7 89.0 74.7 46.6 70.1
 Multi-Modal Reward Models
@@
 Previous-SOTA Previous-SOTA IXC-2.5 IXC-2.5-Chat
-Instruction WildVision(0617) [62] Open 89.2 [31] 67.3 [94] 37.5 74.6
-Following MIA${}_{(\text{val})}$ [68] Open 88.6 [31] 80.7 [90] 80.4 84.0
-& Chat MM-MT${}_{(\text{val})}$ [1] Open 7.72 [31] 5.45 [90] 3.85 5.70
-MM-Vet v2${}_{(\text{0613})}$ [103] Open 71.8 [4] 58.1 [14] 45.8 54.8
-Knowledge MMBench${}_{(\text{v1.1})}$ [52] MCQ 85.7 [74] 82.7 [61] 79.4 79.0
-MMMU${}_{(\text{val})}$ [106] MCQ 70.7 [31] 56.2 [14] 42.9 44.1
-MMStar [12] MCQ 72.7 [74] 63.2 [14] 59.9 59.6
-Reasoning MathVista${}_{(\text{mini})}$ [59] VQA 78.4 [74] 66.5 [60] 63.7 63.4
-MathVerse${}_{(\text{vision-only})}$ [113] VQA 54.8 [26] 26.6 [50] 16.2 19.0
-MathVision${}_{(\text{full})}$ [89] VQA 43.6 [26] 22.0 [50] 17.8 18.8
-Text-Rich TextVQA${}_{(\text{val})}$ [79] VQA 82.0 [64] 78.5 [42] 78.2 81.3
-ChartQA${}_{(\text{test})}$ [63] VQA 81.2 [64] 82.4 [99] 82.2 80.5
-OCRBench [49] VQA 89.4 [74] 82.2 [14] 69.0 70.0
-Figure 3: Visualizations of multi-modal dialogues generated by IXC-2.5-Chat on instruction following abilities.
-Figure 4: Visualizations of multi-modal dialogues generated by IXC-2.5-Chat on open-ended questions.
-Implementation Details
-For IXC-2.5-Reward, the learning rates were set at 1e-5 with a batch size of 256.
-As for IXC-2.5-Chat, the learning rates were set at 5e-5 with a batch size of 256.
-We set the PPO hyper-parameters $\gamma=0.99$, $\beta=0.95$, and $\epsilon=0.2$.
+Instruction WildVision(0617) Lu et al. (2024c) Open 89.2 Hurst et al. (2024) 67.3 Xiong et al. (2024) 37.5 74.6
+Following MIA${}_{(\text{val})}$ Qian et al. (2024) Open 88.6 Hurst et al. (2024) 80.7 Wang et al. (2024d) 80.4 84.0
+& Chat MM-MT${}_{(\text{val})}$ Agrawal et al. (2024) Open 7.72 Hurst et al. (2024) 5.45 Wang et al. (2024d) 3.85 5.70
+MM-Vet v2${}_{(\text{0613})}$ Yu et al. (2023) Open 71.8 Anthropic (2024) 58.1 Chen et al. (2024d) 45.8 54.8
+Knowledge MMBench${}_{(\text{v1.1})}$ Liu et al. (2025) MCQ 85.7 SenseTime (2024) 82.7 Lu et al. (2024b) 79.4 79.0
+MMMU${}_{(\text{val})}$ Yue et al. (2024) MCQ 70.7 Hurst et al. (2024) 56.2 Chen et al. (2024d) 42.9 44.1
+MMStar Chen et al. (2024b) MCQ 72.7 SenseTime (2024) 63.2 Chen et al. (2024d) 59.9 59.6
+Reasoning MathVista${}_{(\text{mini})}$ Lu et al. (2023) VQA 78.4 SenseTime (2024) 66.5 Lu et al. (2024a) 63.7 63.4
+MathVerse${}_{(\text{vision-only})}$ Zhang et al. (2025b) VQA 54.8 Google (2024) 26.6 Liu et al. (2024c) 16.2 19.0
+MathVision${}_{(\text{full})}$ Wang et al. (2024c) VQA 43.6 Google (2024) 22.0 Liu et al. (2024c) 17.8 18.8
+Text-Rich TextVQA${}_{(\text{val})}$ Singh et al. (2019) VQA 82.0 Megvii (2024) 78.5 Li et al. (2024a) 78.2 81.3
+ChartQA${}_{(\text{test})}$ Masry et al. (2022) VQA 81.2 Megvii (2024) 82.4 Yao et al. (2024) 82.2 80.5
+OCRBench Liu et al. (2024d) VQA 89.4 SenseTime (2024) 82.2 Chen et al. (2024d) 69.0 70.0
 5.1 Evaluation Results of IXC-2.5-Reward
-Benchmarks
-To evaluate IXC-2.5-Reward, we use diverse reward model benchmarks: (1) VL-RewardBench [43], encompassing 1250 multi-modal problems addressing general understanding, hallucination, and reasoning challenges; (2) Reward-Bench [41], with 2985 language-only problems including chat, chat hard, safety and reasoning; and (3) RM-Bench [51], comprising 1237 language-only problems across chat, math, code, and safety.
+Benchmarks.
+To evaluate IXC-2.5-Reward, we use diverse reward model benchmarks: (1) VL-RewardBench Li et al.
+(2024b), encompassing 1250 multi-modal problems addressing general understanding, hallucination, and reasoning challenges; (2) Reward-Bench Lambert et al.
+(2024b), with 2985 language-only problems including chat, chat hard, safety and reasoning; and (3) RM-Bench Liu et al.
+(2024b), comprising 1237 language-only problems across chat, math, code, and safety.
 RM-Bench defines three tracks (easy, normal, hard) that evaluate the sensitivity of reward models to subtle content variations and style biases.
@@
 5.1.1 Results on VL-RewardBench
-Main Results
+Main Results.
 Tab.
-3 presents the evaluation results of various multi-modal RMs on the VL-RewardBench [43].
+3 presents the evaluation results of various multi-modal RMs on the VL-RewardBench Li et al.
+(2024b).
 Unlike previous multi-modal generative reward models, our IXC-2.5-Reward is a discriminative model that predicts a scalar reward.
@@
 Notably, IXC-2.5-Reward achieves the highest overall accuracy (65.8%) among open-source models and the highest Macro Accuracy (70.0%) among all models, indicating its superior performance in handling diverse tasks within the VL-RewardBench.
-Strong Performance on General Problems
+Strong Performance on General Problems.
 The results in Table 3 reveal that IXC-2.5-Reward achieves a significantly higher accuracy (84.7%) on general problems compared to other generative RMs.
@@
 5.1.2 Results on Reward Bench and RM-Bench
-Main Results
+Main Results.
 We argue that multi-modal RMs should preserve strong language processing abilities despite the incorporation of image and video data during training.
@@
 The results demonstrate that IXC-2.5-Reward achieves considerable performance and surpasses other multi-modal models on this benchmark.
-Sensitivity to Content and Style
-Consistent with findings in [51], IXC-2.5-Reward demonstrates sensitivity to subtle content variations and style biases, an issue often overlooked in multi-modal research.
+Sensitivity to Content and Style.
+Consistent with findings in Liu et al.
+(2024b), IXC-2.5-Reward demonstrates sensitivity to subtle content variations and style biases, an issue often overlooked in multi-modal research.
 We believe further research is needed to enhance the robustness of multi-modal reward models.
 5.2 Evaluation Results of IXC-2.5-Chat
-Benchmarks
+Benchmarks.
 We select four representative benchmarks for evaluating the instruction following and in-the-wild chatting abilities of LVLMs.
-(1) The WildVision bench [62] uses prompts collected from user submissions, reflecting real-world multimodal interactions.
-(2) MIA-bench [68] that is specially designed to evaluate instruction following.
-(3) MM-MT [1] which is an instruction-following benchmark for multi-modal models, exhibits a strong correlation with LMSys-Vision ELO ratings [15].
-(4) MM-Vet [103] that evaluate LVLMs on complex tasks such as language generation.
-These datasets contain open-ended questions and the referenced answers, and evaluation is performed using an LLM-as-a-Judge [118] approach, which involves using a judge model like GPT-4o [31] to predict scores.
-We also report the performance on other categories, such as MMBench [52], MMMU [106] and MMStar [12] for general knowledge, MathVerse [113] and MathVision [89] for math reasoning, TextVQA [79], ChartQA [63] and OCRbench [49] for text-rich document understanding.
+(1) The WildVision bench Lu et al.
+(2024c) uses prompts collected from user submissions, reflecting real-world multimodal interactions.
+(2) MIA-bench Qian et al.
+(2024) that is specially designed to evaluate instruction following.
+(3) MM-MT Agrawal et al.
+(2024) which is an instruction-following benchmark for multi-modal models, exhibits a strong correlation with LMSys-Vision ELO ratings Chiang et al.
+(2024).
+(4) MM-Vet Yu et al.
+(2023) that evaluate LVLMs on complex tasks such as language generation.
+These datasets contain open-ended questions and the referenced answers, and evaluation is performed using an LLM-as-a-Judge Zheng et al.
+(2023) approach, which involves using a judge model like GPT-4o Hurst et al.
+(2024) to predict scores.
+We also report the performance on other categories, such as MMBench Liu et al.
+(2025), MMMU Yue et al.
+(2024) and MMStar Chen et al.
+(2024b) for general knowledge, MathVerse Zhang et al.
+(2025b) and MathVision Wang et al.
+(2024c) for math reasoning, TextVQA Singh et al.
+(2019), ChartQA Masry et al.
+(2022) and OCRbench Liu et al.
+(2024d) for text-rich document understanding.
 These benchmarks utilize multiple-choice questions (MCQ) or visual question answering (VQA), where responses are limited to short keywords and evaluated based on string matching.
-Results on Instruction Following & Chat
+Results on Instruction Following & Chat.
 Tab.
-6 shows that IXC-2.5-Chat outperforms previous SOTA models across multiple benchmarks (WildVision, MIA, and MM-MT), demonstrating significant improvements in multi-modal understanding with instruction following ability and provide more comprehensive information for in-the-wild chat scenario.
-Results on Other Categories
+6 shows that IXC-2.5-Chat outperforms previous SOTA models across multiple benchmarks (WildVision, MIA, and MM-MT), demonstrating significant improvements in multi-modal understanding with instruction following ability and providing more comprehensive information for in-the-wild chat scenarios.
+Results on Other Categories.
 On other categories (Knowledge, Reasoning, and Text-Rich), IXC-2.5-Chat performs comparably to the supervised fine-tuned (SFT) model IXC-2.5, demonstrating that RL training with IXC-2.5-Reward improves instruction following and conversational ability without sacrificing performance in these areas.
+6 Conclusion and Future Work
+We present IXC-2.5-Reward, a multi-modal reward model that is capable of multi-modal RL training, test-time scaling, and data cleaning.
+Using IXC-2.5-Reward, we further trained IXC-2.5-Chat via RLHF techniques to optimize the multi-modal user chat experience, focusing on providing detailed explanations and in-depth answers.
+We believe that exploring multi-modal reward models with on-policy reinforcement learning algorithms holds significant promise for future research, such as exploring reward benchmarks and RL algorithms for video alignment.
+7 Limitations
+The limitation of our work stems from the composition of our training data, which is primarily sourced from English language corpora.
+This reliance on English-centric data potentially limits the multilingual capabilities of our reward model.
+The English language datasets may reflect specific cultural viewpoints and societal biases prevalent in English-speaking communities.
+Future research should consider the incorporation of multilingual datasets to mitigate these limitations and enhance the generalizability and fairness of the multi-modal reward model.
+Appendix
+Appendix A More Experimental Results
+Implementation Details
+For IXC-2.5-Reward, the learning rates were set at 1e-5 with a batch size of 256.
+As for IXC-2.5-Chat, the learning rates were set at 5e-5 with a batch size of 256.
+We set the PPO hyper-parameters $\gamma=0.99$, $\beta=0.95$, and $\epsilon=0.2$.
 Table 7: Ablation Studies of the impact of response length constraints of reward models that guided training IXC-2.5-Chat.
@@
 As our focus is on optimizing user experience, not benchmark scores, we retain the length constraints.
-Following the precedent set by language-only benchmarks (e.g., [23]), we believe multi-modal Chat Arena and dialogue benchmarks should also address potential length and style biases in their evaluation protocols in future work.
+Following the precedent set by language-only benchmarks (e.g., Dubois et al.
+(2024)), we believe multi-modal Chat Arena and dialogue benchmarks should also address potential length and style biases in their evaluation protocols in future work.
 Results on Test-Time Scaling
@@
 These strengths significantly enhance IXC-2.5-Chat’s effectiveness in multi-modal chat interactions.
-6 Conclusion and Future Work
-We present IXC-2.5-Reward, a multi-modal reward model that is capable of multi-modal RL training, test-time scaling and data cleaning.
-IXC-2.5-Reward supports text, image, and video inputs and is trained on diverse, high-quality preference data to ensure robust preference discrimination.
-Using IXC-2.5-Reward, we further trained IXC-2.5-Chat via RLHF techniques to optimize the multi-modal user chat experience, focusing on providing detailed explanations and in-depth answers.
-We believe that exploring multi-modal reward models with on-policy reinforcement learning algorithms holds significant promise for future research.
-For instance, while IXC-2.5-Reward can score multi-image and video inputs, existing reward benchmarks and applying RL algorithms for video alignment remain scarce.
-Moreover, extending the development of multi-modal reward models beyond chat to other domains, such as process reward models (PRMs) for multi-modal reasoning, is a promising direction.
+Figure 3: Visualizations of multi-modal dialogues generated by IXC-2.5-Chat on instruction following abilities.
+Figure 4: Visualizations of multi-modal dialogues generated by IXC-2.5-Chat on open-ended questions.